OpenStudy (anonymous):

A circle has a radius of 9 ft. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth.

OpenStudy (anonymous):

Hello, i have no clue what to do

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@TheAsker2002

OpenStudy (anonymous):

Im not sure were to begin

OpenStudy (anonymous):

@Legends

TheSmartOne (thesmartone):

Do you know the area formula for a circle?

OpenStudy (anonymous):

No i do not.

TheSmartOne (thesmartone):

Hint: \(\sf\Large Area= 2 \times\pi \times r^2\)

OpenStudy (anonymous):

wait, is it A=πr2

OpenStudy (anonymous):

Ok so what next?

TheSmartOne (thesmartone):

plug in r=9 and pi = 3.14 and just solve.

OpenStudy (anonymous):

28.26?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

2 times 3.14 times 9^2 will never even be close to equalling 28.26

OpenStudy (anonymous):

508.68

TheSmartOne (thesmartone):

correct

OpenStudy (anonymous):

so thats my answe...?

TheSmartOne (thesmartone):

yes

TheSmartOne (thesmartone):

oops @Deadpool117 the correct answer to your question is 254.34 idk where I put in that 2 from :/

TheSmartOne (thesmartone):

@Deadpool117